# DarwinNet: An Evolutionary Network Architecture for Agent-Driven Protocol Synthesis

### 27 Mar 2026 v1

@@ IV-C Protocol Solidification Index (PSI)
Physical Significance and Maturity Measurement: The PSI measures the “collapse” of a system from Slow Thinking (System 2) to Fast Thinking (System 1).
@@
  • Stabilized Phase ($M\to 1$): As the system matures, optimal interaction patterns are “solidified” into efficient bytecode. The Darwin Cortex enters a dormant state, and the network operates at physical limit speeds using its evolved “reflexes.”
This metric reflects the “breathing” nature of DarwinNet: when a network faces a novel attack or environment shift, $M(t)$ temporarily drops as the system “wakes up” to evolve, then rises again as the new solution is codified and optimized [6].
@@ IV-D Safety Guardrails and Protocol Alignment
To mitigate the risks associated with AI-generated logic, DarwinNet implements a multi-stage safety filter. First, the Evolution Engine is restricted by a ”Policy-Aligned Template” that prevents the synthesis of known malicious communication patterns (e.g., amplification attack vectors). Second, the Immune System performs static analysis and formal verification on the synthesized WASM bytecode before execution. Finally, the L0 anchor enforces runtime monitoring, where any protocol mutation that exceeds pre-set traffic or resource quotas is instantly melted and reverted to a known-safe baseline state. This multi-layered defense-in-depth ensures that the self-evolving nature of DarwinNet remains aligned with human intent and network security policies.
@@ V-A Simulation Methodology and Fault Modeling
To simulate “protocol mismatch events” (anomalies where existing protocols fail to meet environmental or intent requirements), we adopt the Crow-AMSAA (Duane) Power Law process [8, 26]. This model is particularly suited for systems undergoing continuous reliability improvement through iterative “discovery and repair” cycles [13].
@@
The occurrence of protocol failure events follows an intensity function in Eq. 1. In our simulation, we set the shape parameter $\beta<1$ (e.g., $\beta=0.6$) to model the evolutionary learning effect. As illustrated in the Duane plot in Fig. 4, the cumulative failure rate exhibits a linear downward trend on a log-log scale. This validates the hypothesis that as DarwinNet adapts to environmental stressors, the frequency of “unseen” edge cases decreases over time, representing the transition from a chaotic state to a governed evolutionary trajectory.
@@ V-B Quantitative Analysis of Protocol Solidification
The primary metric for evaluating DarwinNet’s maturity is the PSI in Eq. 2. Fig. 5 captures the collapse efficiency of the system. Initially, in the “Chaotic Phase,” $M(t)$ is low due to frequent agent interventions. However, as interaction patterns stabilize, $M(t)$ rapidly converges toward 1.0, indicating that the LLM has effectively “compiled” environmental requirements into efficient binary “reflexes.”
@@
To test the system’s robustness, we introduced an environmental shock (a novel, undefined protocol conflict) at $N_{total}=1000$. As shown in Fig. 6, the PSI experience a temporary dip as the Darwin Cortex is re-awakened to handle the anomaly. Crucially, the system exhibits anti-fragility: it doesn’t just recover but rapidly re-solidifies, reaching a new equilibrium [2]. This “V-shaped” recovery demonstrates that DarwinNet treats unexpected errors as fuel for further evolution rather than causes of failure.
@@ V-C Performance Gain and Engineering Feasibility
Figure 7: Latency performance evolution on a logarithmic scale, highlighting the shift from high-latency “Evolutionary Tax” (LLM intervention) to near-native “Efficiency Dividend.”
In the “Initial Evolutionary Period,” the system pays a significant Evolutionary Tax, with latency spikes reaching 500ms during LLM-based reconstruction. However, as the protocol solidifies, the moving average latency drops by three orders of magnitude, eventually aligning with the 1ms physical execution limit of the WASM runtime. This transition into the Efficiency Dividend zone proves that the high cost of agent intervention is a transient investment that yields long-term, near-optimal operational efficiency.
@@ V-D Convergence Verification under Crow-AMSAA Constraints
Finally, by correlating the reliability growth in Fig. 4 with the maturity index in Fig. 5, we verify the mathematical convergence of the framework. The results confirm that under stable environmental constraints, the protocol reaches a Solidification Equilibrium. At this stage, the “growth” process stabilizes, and the protocol is sufficiently adapted to the environment to require no further agent intervention. This validates that DarwinNet is not an infinitely chaotic system but a stable, commercially viable architecture capable of reaching physical performance limits through autonomous growth.
@@ Acknowledgment
The authors contributed collectively to the research. However, the specific strategic visions and conceptual interpretations presented herein are attributed to the corresponding author to reflect his current stage of thinking, and do not necessarily represent the official views of the other authors or their affiliated institutions.

### 13 Apr 2026 v2

@@ IV-C Protocol Solidification Index (PSI)
Physical Significance and Maturity Measurement: The PSI measures the collapse of a system from Slow Thinking (System 2) to Fast Thinking (System 1).
@@
  • Stabilized Phase ($M\to 1$): As the system matures, optimal interaction patterns are solidified into efficient bytecode. The Darwin Cortex enters a dormant state, and the network operates at physical limit speeds using its evolved reflexes.
This metric reflects the breathing nature of DarwinNet: when a network faces a novel attack or environment shift, $M(t)$ temporarily drops as the system wakes up to evolve, then rises again as the new solution is codified and optimized [6].
@@ IV-D Safety Guardrails and Protocol Alignment
To mitigate the risks associated with AI-generated logic, DarwinNet implements a multi-stage safety filter. First, the Evolution Engine is restricted by a Policy-Aligned Template that prevents the synthesis of known malicious communication patterns (e.g., amplification attack vectors). Second, the Immune System performs static analysis and formal verification on the synthesized WASM bytecode before execution. Finally, the L0 anchor enforces runtime monitoring, where any protocol mutation that exceeds pre-set traffic or resource quotas is instantly melted and reverted to a known-safe baseline state. This multi-layered defense-in-depth ensures that the self-evolving nature of DarwinNet remains aligned with human intent and network security policies.
@@ V-A Simulation Methodology and Fault Modeling
To simulate protocol mismatch events (anomalies where existing protocols fail to meet environmental or intent requirements), we adopt the Crow-AMSAA (Duane) Power Law process [8, 26]. This model is particularly suited for systems undergoing continuous reliability improvement through iterative discovery and repair cycles [13].
@@
The occurrence of protocol failure events follows an intensity function in Eq. 1. In our simulation, we set the shape parameter $\beta<1$ (e.g., $\beta=0.6$) to model the evolutionary learning effect. As illustrated in the Duane plot in Fig. 4, the cumulative failure rate exhibits a linear downward trend on a log-log scale. This validates the hypothesis that as DarwinNet adapts to environmental stressors, the frequency of unseen edge cases decreases over time, representing the transition from a chaotic state to a governed evolutionary trajectory.
@@ V-B Quantitative Analysis of Protocol Solidification
The primary metric for evaluating DarwinNet’s maturity is the PSI in Eq. 2. Fig. 5 captures the collapse efficiency of the system. Initially, in the Chaotic Phase, $M(t)$ is low due to frequent agent interventions. However, as interaction patterns stabilize, $M(t)$ rapidly converges toward 1.0, indicating that the LLM has effectively compiled environmental requirements into efficient binary reflexes.
@@
To test the system’s robustness, we introduced an environmental shock (a novel, undefined protocol conflict) at $N_{total}=1000$. As shown in Fig. 6, the PSI experience a temporary dip as the Darwin Cortex is re-awakened to handle the anomaly. Crucially, the system exhibits anti-fragility: it doesn’t just recover but rapidly re-solidifies, reaching a new equilibrium [2]. This V-shaped recovery demonstrates that DarwinNet treats unexpected errors as fuel for further evolution rather than causes of failure.
@@ V-C Performance Gain and Engineering Feasibility
Figure 7: Latency performance evolution on a logarithmic scale, highlighting the shift from high-latency Evolutionary Tax (LLM intervention) to near-native Efficiency Dividend.
In the Initial Evolutionary Period, the system pays a significant Evolutionary Tax, with latency spikes reaching 500ms during LLM-based reconstruction. However, as the protocol solidifies, the moving average latency drops by three orders of magnitude, eventually aligning with the 1ms physical execution limit of the WASM runtime. This transition into the Efficiency Dividend zone proves that the high cost of agent intervention is a transient investment that yields long-term, near-optimal operational efficiency.
@@ V-D Convergence Verification under Crow-AMSAA Constraints
Finally, by correlating the reliability growth in Fig. 4 with the maturity index in Fig. 5, we verify the mathematical convergence of the framework. The results confirm that under stable environmental constraints, the protocol reaches a Solidification Equilibrium. At this stage, the growth process stabilizes, and the protocol is sufficiently adapted to the environment to require no further agent intervention. This validates that DarwinNet is not an infinitely chaotic system but a stable, commercially viable architecture capable of reaching physical performance limits through autonomous growth.
@@ Acknowledgment
The authors contributed collectively to the research. However, the specific strategic visions and conceptual interpretations presented herein are attributed to the corresponding author (xujiinliang@caict.ac.cn, jlxufly@gmail.com) to reflect his current stage of thinking, and do not necessarily represent the official views of the other authors or their affiliated institutions.